OpenStudy (anonymous):

the ratio of 2 numbers is 7 to 5. if the sum of the numbers is 960 what are the numbers

OpenStudy (whpalmer4):

a/b = 7/5 a+b = 960 a = 7b/5 7b/5 + b = 960 7b/5 + 5b/5 = 960 12b = 5*960 b = 5*960/12 = 400 a = 7*400/5 = 560 or: 7 + 5 = 12, 12 parts in 960, each part = 960/12=80 a is 7 parts of the whole, b is 5 parts, so 7*80 = 560 and 5*80 = 400